### v3

@@ 5 AI as a Service
As described in Section 1, to more effectively support “native intelligence” and achieve ubiquitous “universal intelligence”, future 6G networks will treat AIaaS, giving rise to the concept of 6G AIaaS [248], which is presented in Figure 25. 6G will provide efficient end-to-end support for AI-related businesses and applications, intelligently connecting distributed agents for large-scale AI deployment across various industries. 6G AI services will address needs for high real-time performance, high security and privacy, or low overall energy consumption by conducting AI training or inference within the network, offering intelligent capabilities tailored to different application scenarios.
@@ 5.1 Typical Scenarios of AIaaS
Typical scenarios for 6G AIaaS are contexts and situations where AI in 6G networks comes into play, including but not limited to industries and fields such as industry, agriculture, daily life, education, and healthcare [250].
@@ 5.1.1 Immersive Communication
In future communication, immersive technologies like VR, AR, and Measurement Reports will enhance interactive experiences [251]. AIaaS will support these technologies with real-time language translation, emotion analysis, and personalized suggestions in virtual environments, improving communication efficiency. Trained AI models will be validated to ensure realistic and credible interactions. 6G AIaaS will offer comprehensive AI training, collecting user data for multimodal learning and enabling rapid model training and updates [252]. It will support multi-device linkage, allowing various devices to interact with virtual environments. 6G’s extended coverage will seamlessly connect virtual and real worlds, facilitating cross-system information sharing. DT technology will enhance AI validation, generating richer sample data than physical environments and improving model stability and performance. As intelligent terminals increase, 6G’s ultra-large connectivity and low latency will be essential for AI training and validation, supporting precise positioning and massive data collection.
6G AIaaS must deliver high-quality autonomous services to ensure secure interactions in virtual environments, with options ranging from fully autonomous to human-controlled services. This support will lead to more realistic and efficient immersive communication, advancing remote collaboration and communication.
@@ 5.1.2 Intelligent Industrial Robots
Industrial robots will be widely utilized in future industrial manufacturing and transportation scenarios [253]. For instance, construction robots can complete building tasks under remote control. Industrial robots will be considered to replace human workers for hazardous tasks, particularly in extreme working environments such as high temperatures and high altitudes. This will enable an “unmanned factory”, where people only need to use their terminal devices. After completing the AI training of intelligent robots, it is essential to validate the AI models before deployment to ensure the safe and efficient execution of production tasks by the intelligent robots.
6G AIaaS provides AI training services for robots, such as data collection through robots, applying model training for multi-agent learning, and distributing the trained models to the robots. The extreme transmission capabilities of 6G will support rapid model training and parameter exchange between robots and the network. 6G AIaaS also offers multi-system linkage capabilities, allowing various terminal devices to network with robots flexibly and enabling remote control via computers, smartphones, and VR devices. The coverage of the 6G mobile communication system extends from conventional user terminals to environmental sensing devices. Robots can connect and interact with the environment through 6G AIaaS’s sensing technology. They can even engage in cross-environment and cross-system learning with robots from different departments, sharing experiences. Additionally, the 6G network can leverage DT networks to provide AI verification services, iteratively training AI models with greater robustness and better performance through performance pre-validation. 6G AIaaS offers AI verification services for trained models, requiring the DT network to generate sample data in more scenarios than the physical environment, reducing the overhead and performance impact of data collection from the physical network, thereby further improving the stability and performance of AI models.
Intelligent industrial robots demand service quality regarding algorithmic performance metrics, training time, computational precision, duration, and efficiency. With the increasing variety of intelligent terminal types in recent years, 5G may gradually become insufficient to meet the AI training needs of more devices. This necessitates 6G to offer ultra-high connectivity, low latency, and stable AI training services. AI verification services require the 6G network to provide the DT network’s virtual environment with massive data collection, precise positioning, and ultra-low latency transmission capabilities. Furthermore, the intelligent industrial robot scenario requires service quality indicators for the autonomy dimension. The safety of industrial robots can prevent many hazards caused by human negligence or fatigue, requiring 6G AIaaS to provide high-quality services that are fully autonomous, partially human-controlled, and fully human-controlled.
@@ 5.1.3 Precision Medical
In the future, smart healthcare will encompass various aspects of disease prevention, prediction, diagnosis, reasoning, monitoring, clinical surgery, patient care, and vaccine and drug development throughout the entire lifecycle [254]. Smart healthcare will involve centralized hospitals at all levels and distributed community and home healthcare scenarios in the spatial dimension. In the professional dimension, it will integrate and utilize knowledge and experience across different medical disciplines. The overall vision of smart healthcare is to enable individuals to obtain better care, treatment, and protection for their health and well-being at lower costs and more conveniently and efficiently. The new 6G system will better support the massive information transmission and synchronization required for smart healthcare and directly empower the processing and decision-making of medical information.
Healthcare is closely related to individual life and health safety, making the accuracy and reliability of predictive diagnostics crucial. conventional predictive diagnostics based on single AI models face limitations in accuracy, reliability, and real-time performance due to the maturity of AI algorithms and the availability of case data samples. By leveraging AI over 6G networks, geographically dispersed medical institutions and individual practitioners can connect more extensively, aggregate more AI models and case data samples, and rapidly transmit and synchronize information between doctors and patients. This enhances FL and group learning, continually improving the accuracy, reliability, and real-time performance of predictive diagnostics and treatment actions, thereby increasing trust in the prospects of smart healthcare.
Smart healthcare can also achieve more rational planning and optimized allocation of various medical resources through 6G network AI, significantly reducing the physical and psychological stress on doctors and patients and addressing issues such as “difficulties in seeking medical treatment,” “long waiting times”, and “over-medication”. Additionally, protecting the privacy of predictive diagnostics and treatment results for certain diseases is vital. Using a 6G network, AI-distributed technology can better safeguard individual medical information privacy.
@@ 5.1.4 Ultra-Intelligent Transportation
In the future, ultra-intelligent transportation will involve various stages across the temporal dimension, including high-definition map downloads, vehicle environment perception, environmental prediction, and route planning. Spatially, ultra-intelligent transportation will cover scenarios such as highways, urban roads, and enclosed areas. The vision for ultra-intelligent transportation is to enable intelligent vehicles to provide users with a safer, more efficient, and more comfortable riding experience at a lower cost. As the new foundational infrastructure for future ICT, the 6G system boasts ultra-low latency, high-reliability transmission, supercomputing power, and precise sensing. These characteristics will support the real-time transmission, processing, and intelligent decision-making of data related to ultra-intelligent transportation, achieving full network intelligence.
In the context of ultra-intelligent transportation, taking autonomous driving as an example, the 6G system can provide AI-based data services, AI computation offloading services, and AI-based environment prediction or path planning. AI-based data services refer to the 6G system providing autonomous vehicles with AI-driven environmental perception results. For instance, when an autonomous vehicle encounters a perception blind spot, the 6G system can collect perception data through sensors or wireless signals, then use AI models to infer the environment within the blind spot and feed the results back to the vehicle, enhancing driving safety. AI computation offloading refers to the process where autonomous vehicles offload part of their AI model inference or training computations to the 6G system. For example, when a vehicle drives through a complex road segment, the computational power required for AI model inference can increase significantly, resulting in unacceptable latency. In such cases, the vehicle can offload some of the AI computation tasks to the 6G system, thereby reducing the processing burden on the vehicle’s chips.
Targeting sea, land, air, and space transportation, the 6G network’s extensive coverage and intelligent capabilities enable autonomous driving and smart traffic management for vehicles, intelligent scheduling for ships and aviation, and unified control and management of water, land, underground, and aerial transportation. With the 6G network’s DL and multi-sensor fusion capabilities, an integrated perception and decision-making architecture is established for vehicles, roadside infrastructure, and cloud platforms. This architecture advances a new era of intelligent transportation characterized by intelligent travel, ubiquitous services, and comprehensive management and control. In terms of travel, it achieves the organic unification of individual optimal travel solutions and the overall highest efficiency of the transportation system.
@@ 5.2 QoAIS
QoAIS refers to a system of metrics framework for AI services in the network, which broadly needs to involve dimensions such as performance, connectivity, computation, data, security, and orchestration [255]. The performance dimension metrics in QoAIS differ from the model KPI for ML. This is because the latter refers to the evaluation metrics of the model itself, which are mainly used for optimization and tuning of the model. The former extends that content, such as generalizability and robustness of the model/algorithm, to describe the whole system more comprehensively in terms of performance. The metrics of other dimensions in QoAIS are new metrics that consider the 6G network itself and thus extend it. As shown in Table 3, the content of the whole QoAIS system should include:
@@
Performance-related: boundary of performance metrics, training time, generalization, reusability, robustness, interpretability, consistency between loss function and optimization objectives, fairness, etc. 1) Boundary of Performance Metrics: Defines the upper and lower limits for assessing the quality of model performance indicators, such as error rate, precision, recall, etc. 2) Generalization: The ability of a model, after training, to accurately predict outcomes on new, unseen data. 3) Reusability: The capability of a model to continue functioning effectively when applied to varying scenarios. 4) Robustness: The characteristic of a model to maintain certain performance levels despite perturbations, attacks, or uncertainty in the input data. 5) Interpretability: The extent to which a model supports understanding its internal mechanisms and the outcomes it produces. 6) Consistency Between Loss Function and Optimization Objectives: The alignment between the design of the loss function during model training and the optimization goals specific to AI use cases, including whether all variables relevant to the optimization targets are fully accounted for in the function.
Connection-related: bandwidth and jitter, link delay and jitter, bit error rate and jitter, reliability, etc.
Data-related: feature redundancy, completeness, data accuracy, time-consuming data preparation, sample space balance, sample distribution dynamics, etc.
Computation-related: computing accuracy, duration, efficiency, etc.
Security-related: information confidentiality, data/algorithm privacy levels, data authenticity, traceability, etc.
Orchestration-related: full autonomy, partially controllable by humans, fully controllable by humans, etc. Autonomy refers to the requirements for autonomous operation and human intervention in the AI data/training/validation/inference service workflow, reflecting users’ expectations for the degree of automation in AI services. It is classified into three levels: 1) Full Autonomy: Fully automated AI services requiring no human intervention throughout the process. 2) Partially Controllable by Humans: AI service workflows that are automated in some stages but require human assistance in others. 3) Fully Controllable by Humans: AI service workflows necessitate human involvement in all stages.
@@ 5.3 Key Capabilities for Supporting AIaaS
The realization of AIaaS essentially means that the entire network serves as a platform capable of providing AI services for users and network devices, which requires the support of many key capabilities. This section will enumerate these key capabilities and potential technologies from the following three aspects.
1) Interconnection of User-Data-Model: To construct an AIaaS platform, it is essential to have extensive data support. This can be achieved by leveraging the network’s inherent sensing capabilities, along with the data collection abilities of numerous user terminals and IoT devices, to enable real-time data acquisition and storage. Furthermore, while ensuring data privacy, it is crucial to establish a direct connection between users and data. This approach provides essential data resources for model training and inference services within the AIaaS platform. Secondly, AI services rely on AI models, and different services require the design of appropriate algorithms to implement intelligent data processing and decision-making processes. Users can utilize the model resources provided by the network or customize models as needed. Therefore, a dedicated interface should be established to connect users with models, providing necessary support for customizing AI models.
2) Management of Heterogeneous Resources and Task Scheduling: 6G networks feature ubiquitous distributed computing power, with CPU/GPU capabilities and cache resources spread across various CN nodes, BS, user devices, and third-party equipment. Using software defined network, computing resources at different nodes can be discovered and tagged, and resources can be uniformly measured through virtualization. Further employing distributed resource management protocols, such as distributed hash tables, enables efficient storage and retrieval of computing resource information. This approach also facilitates automatic updates and timely processing of resource information. Under the premise of resource visibility, the rational allocation of computing, caching, and bandwidth resources significantly impacts the performance of AI services. Techniques such as DRL algorithms, model predictive control, game theory, and dynamic caching can be employed to achieve adaptive resource allocation and service scheduling based on priority or fairness. These methods ensure optimal service performance across varying service loads.
3) Security Assurance in the AIaaS Lifecycle: With the ubiquitous distribution of cloud-edge computing, the demand for interconnecting computing nodes has intensified, complicating the security landscape of network connections. The dynamic scheduling of computing power introduces new requirements for point-to-point network connections and dynamic connection mechanisms across systems, domains, and even borders. To ensure the security of AI services throughout their lifecycle, comprehensive protection of data security, computational security, model security, and privacy is essential. To address these challenges, the constructed AIaaS platform should implement secure data acquisition protocols and device authentication mechanisms to ensure the legitimacy and integrity of data sources. This involves employing secure data fusion algorithms to validate and cleanse data from various terminal devices. Techniques such as homomorphic encryption can be utilized to perform aggregation operations without compromising data privacy. For large-scale data storage, effective storage and timely processing are essential. Techniques like differential privacy and FL should be employed to defend against model inversion and membership inference attacks, thereby safeguarding model parameters and training data security. Additionally, in AI services, the trained models possess significant value and intellectual property. The AIaaS platform should implement mechanisms to protect these models from unauthorized access and misuse. This includes dynamic protection strategies, such as monitoring data flow and model outputs, to safeguard against illegal activities.
In the practical application of AIaaS platforms, requirements such as data management, computational scheduling, and security protection are not isolated; they must be considered across multiple dimensions. These platforms should provide a multidimensional service experience tailored to user-specific needs. The advent of 6G networks will transcend convetional pipeline limitations by offering AIaaS services through an inherently intelligent network architecture. This will address the digital and intelligent needs of the network itself, various industries, and consumers, truly enabling the era of intelligent inclusivity.

### v4

@@ 5 AI as a Service
To more effectively support “native intelligence” and achieve ubiquitously “universal intelligence”, future 6G networks will treat AIaaS, giving rise to the concept of 6G AIaaS [271], which is presented in Figure 22. 6G will provide efficient end-to-end support for AI-related businesses and applications, intelligently connecting distributed agents for large-scale AI deployment across various industries. 6G AI services will address needs for high real-time performance, high security, and privacy, or low overall energy consumption by conducting AI training or inference within the network, offering intelligent capabilities tailored to different application scenarios.
@@ 5.1 Typical Scenarios of AIaaS
Typical scenarios for 6G AIaaS are contexts and situations where AI in 6G networks comes into play, including but not limited to industries and fields such as agriculture, education, and healthcare [131].
@@ 5.1.1 Immersive Communication
In future communication, immersive technologies like VR, AR, and Measurement Reports will enhance interactive experiences [273]. AIaaS will support these technologies with real-time language translation, emotion analysis, and personalized suggestions in virtual environments. Trained AI models will be validated to ensure realistic and credible interactions. 6G AIaaS will offer comprehensive AI training, collecting user data for multimodal learning and enabling rapid model training and updates [274]. It will support multi-device linkage, allowing various devices to interact with virtual environments. 6G’s extended coverage will seamlessly connect virtual and real worlds, facilitating cross-system information sharing. DT technology will enhance AI validation, generating richer sample data than physical environments and improving model stability and performance. As the number of intelligent terminals increases, 6G’s ultra-large connectivity and low latency will be essential for AI training and validation, supporting precise positioning and massive data collection. 6G AIaaS must deliver high-quality autonomous services to ensure secure interactions in virtual environments, with options ranging from fully autonomous to human-controlled services. This support will lead to more realistic and efficient immersive communication, advancing remote collaboration and communication.
@@ 5.1.2 Intelligent Industrial Robots
Industrial robots will be widely utilized in future industrial manufacturing and transportation scenarios [275]. 6G AIaaS provides AI training services for robots, such as data collection through robots, applying model training for multi-agent learning, and distributing the trained models to the robots. The extreme transmission capabilities of 6G will support rapid model training and parameter exchange between robots and the network. 6G AIaaS also offers multi-system linkage capabilities, allowing various terminal devices to network with robots flexibly and enabling remote control via computers, smartphones, and VR devices [276]. Robots can connect and interact with the environment through 6G AIaaS’s sensing technology. They can even engage in cross-environment and cross-system learning with robots from different departments, sharing experiences. Additionally, the 6G network can leverage DT networks to provide AI verification services, iteratively training AI models with greater robustness and better performance through performance pre-validation. 6G AIaaS offers AI verification services for trained models, requiring the DT network to generate sample data in more scenarios than the physical environment, reducing the overhead and performance impact of data collection from the physical network.
@@ 5.1.3 Precision Medical
Smart healthcare will encompass various aspects of disease prevention, prediction, diagnosis, reasoning, monitoring, clinical surgery, patient care, and vaccine and drug development throughout the entire lifecycle [277]. The new 6G system will better support the massive information transmission and synchronization required for smart healthcare and directly empower the processing and decision-making of medical information. By leveraging AI over 6G networks, geographically dispersed medical institutions, and individual practitioners can connect more extensively, aggregate more AI models and case data samples, and rapidly transmit and synchronize information between doctors and patients. This enhances FL and group learning, continually improving the accuracy, reliability, and real-time performance of predictive diagnostics and treatment actions. Smart healthcare can also achieve more rational planning and optimized allocation of various medical resources through 6G network AI, significantly reducing the physical and psychological stress on healthcare practitioners and patients.
@@ 5.1.4 Ultra-Intelligent Transportation
Ultra-intelligent transportation will involve various stages across the temporal dimension, including high-definition map downloads, vehicle environment perception, environmental prediction, and route planning [278]. Taking autonomous driving as an example, the 6G system can provide AI-based data services, AI computation offloading services, and AI-based environment prediction or path planning. AI-based data services refer to the 6G system providing autonomous vehicles with AI-driven environmental perception results. For instance, when an autonomous vehicle encounters a perception blind spot, the 6G system can collect perception data through sensors or wireless signals, then use AI models to infer the environment within the blind spot and feed the results back to the vehicle, enhancing driving safety. AI computation offloading refers to the process where autonomous vehicles offload part of their AI model inference or training computations to the 6G system. For example, when a vehicle drives through a complex road segment, the computational power required for AI model inference can increase significantly, resulting in unacceptable latency. In such cases, the vehicle can offload some of the AI computation tasks to the 6G system, thereby reducing the processing burden on the vehicle’s chips.
With the 6G network’s DL and multi-sensor fusion capabilities, an integrated perception and decision-making architecture is established for vehicles, roadside infrastructure, and cloud platforms [279]. This architecture advances a new era of intelligent transportation characterized by intelligent travel, ubiquitous services, and comprehensive management and control.
@@ 5.2 QoAIS
Quality of AI service (QoAIS) refers to a comprehensive metrics framework for evaluating AI services within a network. It encompasses key dimensions such as performance, connectivity, computation, data, security, and orchestration [280, 272, 39, 271]. Notably, the performance metrics in QoAIS differ from traditional model KPIs used in ML. While model KPIs primarily assess the internal evaluation metrics of a model—serving purposes like optimization and fine-tuning—QoAIS extends this scope. It incorporates factors such as the generalizability and robustness of models or algorithms, thereby offering a holistic view of system-wide performance. The metrics for other dimensions within QoAIS are novel, as they specifically account for the unique characteristics and requirements of 6G networks. As shown in Table 3, the QoAIS system should include:
@@
Performance-related indicators. Aspects critical to evaluating and improving AI models are encompassed, including the boundary of performance metrics, training time, generalization, reusability, robustness, interpretability, and consistency between the loss function and optimization objectives. Boundary of performance metrics defines the upper and lower limits for assessing the quality of model performance indicators, such as error rate, precision, and recall. Generalization reflects a model’s ability to make accurate predictions on new, unseen data, while reusability highlights its capability to remain effective across diverse scenarios. Robustness ensures the model maintains consistent performance despite perturbations, adversarial attacks, or input uncertainties. Interpretability focuses on the extent to which the model’s internal workings and outputs can be understood and explained, fostering transparency and trust. Consistency between the loss function and optimization objectives ensures alignment between the design of the loss function during training and the AI system’s goals, accounting for all relevant variables to achieve the desired outcomes.
Connection-related indicators. Reliable and efficient AI services depend heavily on network characteristics, including bandwidth, jitter, link delay, bit error rate, and overall connection reliability. Bandwidth represents the data transmission capacity of the network, directly impacting the speed and quality of real-time AI applications, such as video analysis or interactive systems. Jitter refers to the variability in packet arrival times, which can lead to inconsistencies in service performance, particularly in latency-sensitive tasks. Link delay quantifies the time taken for data to travel across the network, with lower delays contributing to smoother and more responsive AI systems. Bit error rate measures the rate at which errors occur during data transmission, emphasizing the importance of accurate and reliable data exchange. These factors collectively define the stability and dependability of the connection, directly influencing the QoAISs.
Data-related indicators. The quality and characteristics of data play a vital role in ensuring the success of AI systems, focusing on feature redundancy, completeness, accuracy, and preparation time. Feature redundancy highlights the presence of overlapping or irrelevant features, which can complicate model training and reduce efficiency. Completeness ensures that all necessary data attributes are available to achieve comprehensive insights and predictions. Data accuracy emphasizes the importance of reliable, error-free data to maintain the validity of AI outputs. Time-consuming data preparation reflects the challenges involved in cleaning, transforming, and organizing data for training purposes. Sample space balance ensures equitable representation of different classes or categories within the dataset, critical for minimizing bias. Sample distribution dynamics track changes in data distributions over time, enabling models to adapt and remain effective in evolving environments. Together, these aspects underpin the foundation for building robust and trustworthy AI models.
Computation-related indicators. Efficiency and accuracy in computation are crucial for ensuring reliable and high-performing AI systems, encompassing factors such as computing accuracy, duration, and efficiency. Computing accuracy reflects the precision of the computational results, ensuring the reliability of outputs across tasks. This dimension emphasizes the efficiency and accuracy of AI computations, which include computing accuracy, duration, and efficiency. Duration measures the time taken to complete computational processes, with faster computations enabling real-time or near-real-time performance in AI systems. Efficiency evaluates the resource utilization of the computation, including memory, processing power, and energy consumption. Optimizing these aspects ensures that AI services are not only accurate but also cost-effective and sustainable.
Security-related indicators. Protecting data, algorithms, and systems is a fundamental requirement for AI services, with key considerations including confidentiality, privacy, authenticity, and traceability. Information confidentiality ensures that sensitive data remains inaccessible to unauthorized entities, safeguarding user trust and compliance with regulations. Data/algorithm privacy levels highlight the measures taken to protect proprietary algorithms and personal data, reducing the risk of exposure or misuse. Data authenticity guarantees that the data used in AI processes is genuine and unaltered, ensuring reliable outputs. Traceability refers to the ability to track the origin, flow, and modifications of data and algorithms, enabling accountability and transparency in AI systems. Collectively, these factors contribute to the safe and ethical deployment of AI services.
Orchestration-related indicators. The level of automation and human involvement in AI services plays a key role in determining how workflows are managed and controlled. Autonomy in AI services defines the degree of automation and the extent of human involvement required across data, training, validation, and inference workflows. It reflects user expectations regarding the level of automation and is categorized into three levels. At the highest level, full autonomy enables AI services to operate independently without human intervention. In contrast, partial human control involves workflows where some stages are automated, but others require human assistance. Finally, full human control necessitates human involvement at every stage of the workflow, with no reliance on automation.
@@ 5.3 Key Capabilities for Supporting AIaaS
The concept of AIaaS envisions a network-wide platform that seamlessly delivers AI services to users and devices connected. Achieving this vision hinges on several critical capabilities, which can be grouped into three primary aspects:
5.3.1 Data Support and Model Accessibility
A fundamental requirement for an AIaaS platform is robust data support. This can be achieved by utilizing the network’s sensing capabilities and the data collection functions of IoT devices and user terminals to enable real-time data acquisition and storage. Ensuring data privacy while maintaining a direct link between users and their data is crucial, as this connection underpins model training and inference services. Moreover, AI services also rely on diverse models tailored to different tasks. The platform should allow users to access pre-existing models or create custom models to suit their needs. A dedicated interface connecting users to these models would facilitate customization and enhance the flexibility of AI services.
5.3.2 Distributed Computing and Resource Optimization
The distributed computing power inherent in 6G networks—spanning CPUs, GPUs, and cache resources across CN nodes, BSs, user devices, and third-party systems—forms the backbone of AIaaS. By leveraging software defined network, these resources can be discovered, tagged, and virtualized for unified management [281]. Protocols like distributed hash tables can be employed to store, retrieve, and update resource information efficiently. With resource visibility in place, computing, caching, and bandwidth resources can be allocated dynamically to meet service demands. Techniques such as DRL, game theory, and model predictive control enable adaptive resource allocation and service scheduling, ensuring efficient operation even under varying workloads.
5.3.3 Security and Lifecycle Safeguards
The distribution of cloud-edge computing nodes introduces new challenges in securing dynamic and cross-domain network connections. Comprehensive security measures are needed to protect AI services throughout their lifecycle. These include protocols for secure data acquisition and device authentication to ensure the integrity of inputs. Secure data fusion algorithms can validate and clean data from multiple devices, while homomorphic encryption offers privacy-preserving data aggregation. Safeguarding models, which represent significant intellectual property, is equally important. Measures like access control, differential privacy, FL, and real-time monitoring of data flow can protect against unauthorized use, model inversion, and other emerging threats.
Integrating these capabilities into AIaaS platforms requires a cohesive approach that addresses data management, computational scheduling, and security concerns simultaneously. The multidimensional service experience provided by these platforms should align with user-specific needs. With the evolution of 6G networks, traditional network limitations will be overcome, enabling AIaaS to thrive within an intelligent, adaptive, and efficient network architecture.